OpenStudy (anonymous):

What is the equation of the circle with center (2, –5) that passes through the point (–2, 10)?

OpenStudy (anonymous):

The choices are : (x – 2)2 + (y – (–5))2 = 25 (x – (–2))2 + (y – 10)2 = 241 (x – 2)2 + (y – (–5))2 = 241 (x – (–2))2 + (y – 10)2 = 25

OpenStudy (anonymous):

Distance between center and point on circle = radius. Find that first. Then, equation is: (x - x1)^2 + (y - y1)^2 = r^2

OpenStudy (anonymous):

Where x1, y1 is the center of circle.

OpenStudy (anonymous):

(x – 2)2 + (y – (–5))2 = 241